Please i need your help in solving this equation 2 - 3 ( a + 5 ) = -10
Answer by harpazo(655)

2-3(a+5)= -10
2-3a-15= -10
-3a-13 = -10
-3a = -10+13
-3a = 3
a = 3/-3
a = -1